OpenStudy (anonymous):

what is one of the solutions to the following system of equations y^2+x^2=53 y-x=5

OpenStudy (anonymous):

rewrite the question please

OpenStudy (anonymous):

y^2 +x^2=53 and y-x=5

OpenStudy (anonymous):

what is one of the solutions to the following system of equations

OpenStudy (anonymous):

@mayafrancis

OpenStudy (whpalmer4):

\[y^2+x^2=53\]\[y-x=5\] To solve this, solve the second equation for either \(x\) or \(y\). Then substitute that expression in place of the variable you solved for in the first equation. Now you should have an equation in terms of just one variable which you can solve. When you've done that, plug the value(s) found back into the second equation to find the value of the other variable.

OpenStudy (anonymous):

So like y=5+x?

OpenStudy (triciaal):

yes. now put that in the first first eqn

OpenStudy (triciaal):

where you have y replace it with (5 + x)

OpenStudy (triciaal):

do you need more help? did you get 2, -7

OpenStudy (whpalmer4):

\(y=5+x\) would be a fine start. then in the first equation, wherever you see \(y\), replace it with \((5+x)\) and solve.

OpenStudy (whpalmer4):

when you've solved that equation for \(x\) (there will be two solutions, because it has \(x^2\) as the highest power of \(x\)), take your solutions and use them in \(y = 5+x\) to find the corresponding values of \(y\).

OpenStudy (anonymous):

y=5+x +x^2 = 53?

OpenStudy (triciaal):

y^2 + x^2 =53 (5+x)^2 +x^2=53

OpenStudy (anonymous):

Hmm idk how to solve that

OpenStudy (triciaal):

yes you do i'll work with you

OpenStudy (whpalmer4):

First step is to expand it out: \[(x+5)^2 + x^2 = 53\]\[(x+5)(x+5) + x^2 = 53\] What do you get when you do the multiplication?

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

x2+10x+25

OpenStudy (triciaal):

now add the x^2 and set = 53

OpenStudy (triciaal):

what do you have

OpenStudy (anonymous):

wait add x2 to wat

OpenStudy (triciaal):

look at the original Y^2 + x^2 = 53 then you substituted for y so you get (x+5)^2 + x^2 = 53 you expanded (x+5)^2 to get X^2 + 10x + 25 so now you have x^2 +10x +25 + x^2 = 53

OpenStudy (triciaal):

now group like terms and set = 0

OpenStudy (anonymous):

2x^2+10x+25=53

OpenStudy (anonymous):

and then minus 25?

OpenStudy (anonymous):

can i plz have the answer lol i gotta get this done

OpenStudy (whpalmer4):

\[2x^2+10x + 25 = 53\]Subtract 53 from both sides to get in quadratic formula arrangement.\[2x^2+10x+25-53=0\]\[2x^2+10x-28 = 0\]\[a=2,b=10,c=-28\]\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

-124?

OpenStudy (anonymous):

helloooo

OpenStudy (triciaal):

2X^2 + 10X-28 =0 DIVIDE BY 2 X^2 + 5X-14 = 0 2(X+7)(X-2) =0 X=-7 X=2 you have x values find the corresponding y remember y= x +5 when x = -7, y = -7+5 = -2 when x = 2, y = 2 + 5 = 7

OpenStudy (triciaal):

one solution (2 ,7)

OpenStudy (triciaal):

to check in eqn 1 left side x^2 + y^2 = 2^2 + 7^2 = 4 + 49 = 53 right side 53 53 = 53 x=2, y = 7 is a solution

OpenStudy (anonymous):

Thank you so much ::)

OpenStudy (triciaal):

welcome

OpenStudy (whpalmer4):

\[a=2, b= 10, c = -28\]\[x=\frac{-b\pm\sqrt{b^2 - 4 a c}}{2a} = \frac{-10\pm\sqrt{10^2-4*2*(-28)}}{2*2} \]\[=\frac{1}{4}*(-10\pm\sqrt{100+224})\]\[=\frac{1}4*(-10\pm\sqrt{324}) = \frac{1}{4}(-10\pm18)\]\[x=2 \rightarrow y = x+5 = 2+5 = 7 \rightarrow (2,7)\text { is a solution}\]\[x=-7\rightarrow y = -7+5=-2\rightarrow(-7, -2)\text{ is a solution}\] For this equation, factoring is probably faster for most. For a less friendly one, or one which couldn't be factored, the quadratic formula becomes more appealing.